OpenStudy (anonymous):

Be f:R->R derivable up to the secound derivative of that f''-f=0 prove that there is a constant A for all x so that [(f-Ae^(-x))/(e^x)]'=0

OpenStudy (anonymous):

\[\left[ \frac{ f(x) - Ae ^{-x} }{ e ^{x} } \right]'=0\]

OpenStudy (primeralph):

Well, for f'-f=0, solving the DE, f = Be^(x). So, (Be^(x)-Ae^(-x))/e^x = B- Ae^(-2x) The differential is 2Ae^(-2x) = 0. For this to be true for all x, A = 0.

OpenStudy (anonymous):

f '' - f=0

OpenStudy (primeralph):

Just use the same approach.

OpenStudy (anonymous):

why from: B-Ae^(-2x) u get 2Ae^(-2x)=0 ????

OpenStudy (primeralph):

Because I divided by e^x. Just start it again.

OpenStudy (anonymous):

no u didnt understand!!....

OpenStudy (anonymous):

I see u have B-Ae^(2x)

OpenStudy (primeralph):

I differentiated.

OpenStudy (anonymous):

for f '' - f=0 the only difference is that is B^2 instead of B right?

OpenStudy (primeralph):

No. B is an arbitrary constant.

OpenStudy (anonymous):

but if f(x)=e^x if u derivate u get e^x but if u derivate and get Be^x the primitive is e^cx right?

OpenStudy (primeralph):

f '' - f=0 General solution is Be^(-x) + Be^x

OpenStudy (primeralph):

@matheusoliveira Continue with what I just told you.

OpenStudy (anonymous):

its Be^(-x) + Ce^x the constants need to be different

OpenStudy (anonymous):

how so? i dont think they need to...

OpenStudy (primeralph):

Sorry, they are different.

OpenStudy (anonymous):

so we get [(Be^x + Ce^-x -Ae^-x)/e^x]'=2Ae^(-2x) -2Ce^(-2x)=0

OpenStudy (anonymous):

A/e^2x=C/e^2x so A=C right?

OpenStudy (primeralph):

Yeah I guess